A baker purchased 12lb. of wheat flour and 15lb. of rye flour for a total cost of 18.30. A second purchase, at the same prices, included 15lb. of wheat flour and 10lb. of rye flour. the cost of the second purchase was 16.75. Find the cost per pound of wheat and rye flours.
Answer by Paul(988)

Let the wheat flour be x
Let the rye flour be y
First equatioN:

Second equation:

Solve for y in the second equation:
10y=16.75-15x
y=1.675-1.5x (subsitution)
Subsitute the value in the first equation:
12x+15(1.675-1.5x)=18.3
12x-22.5x=18.3-25.125
-10.5x=-6.825
x=0.6825




Hence the cost of the wheat flour is $0.68 and the rye flour cost $0.65
Paul.